# Make Your Home Safe: Time-aware Unsupervised User Behavior Anomaly Detection in Smart Homes via Loss-guided Mask

### v1

@@ Abstract.
Smart homes, powered by the Internet of Things, offer great convenience but also pose security concerns due to abnormal behaviors, such as improper operations of users and potential attacks from malicious attackers. Several behavior modeling methods have been proposed to identify abnormal behaviors and mitigate potential risks. However, their performance often falls short because they do not effectively learn less frequent behaviors, consider temporal context, or account for the impact of noise in human behaviors. In this paper, we propose SmartGuard, an autoencoder-based unsupervised user behavior anomaly detection framework. First, we design a Loss-guided Dynamic Mask Strategy (LDMS) to encourage the model to learn less frequent behaviors, which are often overlooked during learning. Second, we propose a Three-level Time-aware Position Embedding (TTPE) to incorporate temporal information into positional embedding to detect temporal context anomaly. Third, we propose a Noise-aware Weighted Reconstruction Loss (NWRL) that assigns different weights for routine behaviors and noise behaviors to mitigate the interference of noise behaviors during inference. Comprehensive experiments on three datasets with ten types of anomaly behaviors demonstrates that SmartGuard consistently outperforms state-of-the-art baselines and also offers highly interpretable results.
User Behavior Modeling, Anomaly Detection, Transformer.
copyright: acmlicensedjournalyear: 2024doi: 10.1145/3637528.3671708conference: the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining ; July 14–18, 2024; Barcelona, Spainisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Human and societal aspects of security and privacy
@@ 1. Introduction
Deep learning has been employed across various domains to mine correlations between behaviors for modeling user behavior sequences (Tang et al., 2022, 2023; Li et al., 2024). DeepMove(Feng et al., 2018) leverages RNNs to model both long and short-term mobility patterns of users for human mobility prediction. To capture the dynamics of user’s behaviors, SASRec (Kang and McAuley, 2018) proposes a self-attention based model to achieve sequential recommendation. More recent efforts(Chen et al., 2019; Sun et al., 2019; de Souza Pereira Moreira et al., 2021) primarily focus on transformer-based models for their superior ability to handle sequential behavior data.
@@ Acknowledgements.
We thank the anonymous reviewers for their constructive feedback and comments. This work is supported by the Major Key Project of PCL under grant No. PCL2023A06-4, the National Key Research and Development Program of China under grant No. 2022YFB3105000, and the Shenzhen Key Lab of Software Defined Networking under grant No. ZDSYS20140509172959989.

### v2

@@ Abstract.
Smart homes, powered by the Internet of Things, offer great convenience but also pose security concerns due to abnormal behaviors, such as improper operations of users and potential attacks from malicious attackers. Several behavior modeling methods have been proposed to identify abnormal behaviors and mitigate potential risks. However, their performance often falls short because they do not effectively learn less frequent behaviors, consider temporal context, or account for the impact of noise in human behaviors. In this paper, we propose SmartGuard, an autoencoder-based unsupervised user behavior anomaly detection framework. First, we design a Loss-guided Dynamic Mask Strategy (LDMS) to encourage the model to learn less frequent behaviors, which are often overlooked during learning. Second, we propose a Three-level Time-aware Position Embedding (TTPE) to incorporate temporal information into positional embedding to detect temporal context anomaly. Third, we propose a Noise-aware Weighted Reconstruction Loss (NWRL) that assigns different weights for routine behaviors and noise behaviors to mitigate the interference of noise behaviors during inference. Comprehensive experiments demonstrate that SmartGuard consistently outperforms state-of-the-art baselines and also offers highly interpretable results.
User Behavior Modeling, Anomaly Detection, Transformer, Smart Homes
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671708isbn: 979-8-4007-0490-1/24/08ccs: Security and privacy Human and societal aspects of security and privacy
@@ 1. Introduction
Deep learning has been employed across various domains to mine correlations between behaviors for modeling user behavior sequences (Tang et al., 2022, 2023; Li et al., 2024) and address security issues (Gao et al., 2023, 2024b, 2024a; Bai et al., 2024). DeepMove(Feng et al., 2018) leverages RNNs to model both long and short-term mobility patterns of users for human mobility prediction. To capture the dynamics of user’s behaviors, SASRec (Kang and McAuley, 2018) proposes a self-attention based model to achieve sequential recommendation. More recent efforts(Chen et al., 2019; Sun et al., 2019; de Souza Pereira Moreira et al., 2021) primarily focus on transformer-based models for their superior ability to handle sequential behavior data.
@@ Acknowledgements.
We thank the anonymous reviewers for their constructive feedback and comments. This work is supported by the Major Key Project of PCL under grant No. PCL2023A06-4, the National Key Research and Development Program of China under grant No. 2022YFB3105000, and the Shenzhen Key Lab of Software Defined Networking under grant No. ZDSYS20140509172959989. The first author, Jingyu Xiao, in particular, wants to thank his parents Yingchao Xiao, Aiping Li and his girlfriend Liudi Shen for their kind support. He also thanks for all his friends who are the antidote during his tired period.